A telephone company's goal is to have no more than five monthly line failures on any 100 miles of line. The company currently experiences an average of 4 monthly line failures per 50 miles of line. Let x denote the number of monthly line failures per 100 miles of line. Assuming x has a Poisson distribution:

(a) Find the probability that the company will meet its goal on a particular 100 miles of line.

(b) Find the probability that the company will not meet its goal on a particular 100 miles of line.

(c) Find the probability that the company will have no more than five monthly failures on a particular 200 miles of line.

(d) Find the probability that the company will have more than 12 monthly failures on a particular 150 miles of line.
Answer #1

x follows poisson distribution

\lambda=4/50miles

a)\lambda=2*4=8/100 miles

P(k\leq 5)=\sum_{k=0}^5\frac{e^{-8}8^k}{k!} =0.1912

b)P(k\geq 5)=1-P(k\leq 5)=1-\sum_{k=0}^5\frac{e^{-8}8^k}{k!} =1-0.1912=0.8088

c)\lambda=4*4=16/200 miles

P(k\leq 5)=\sum_{k=0}^5\frac{e^{-16}16^k}{k!}=0.0014

d)\lambda=4*3=12/150miles

P(k>12)=1-\sum_{k=0}^{12}\frac{12^ke^{-12}}{k!}=1-0.576=0.424